# Efficient Splitting Heads

### v1

@@ Abstract
Vision foundation models (VFMs) have demonstrated remarkable performance across a wide range of downstream tasks. While several VFM adapters have shown promising results by leveraging the prior knowledge of VFMs, we identify two inefficiencies in these approaches. First, the interaction between convolutional neural network (CNN) and VFM backbone triggers early layer gradient backpropagation. Second, existing methods require tuning all components, adding complexity. Besides, these adapters alter VFM features, underutilizing the prior knowledge. To tackle these challenges, we propose a new approach called ViT-Split, based on a key observation: the layers of several VFMs, like DINOv2, can be divided into two distinct components: an extractor for learning low-level features and an adapter for learning task-specific features. Leveraging this insight, we eliminate the CNN branch and introduce two heads, task head and prior head, to the frozen VFM. The task head is designed to learn task-specific features, mitigating the early gradient propagation issue. The prior head is used to leverage the multi-scale prior features from the frozen VFM, reducing tuning parameters and overfitting. Extensive experiments on various tasks (e.g., segmentation, detection, depth estimation, and visual question answering) validate the effectiveness and efficiency of ViT-Split. Specifically, ViT-Split reduces training time up to 4×4$\times$4 × while achieving comparable or even better results on ADE20K, compared to other VFM adapters. The project is available: https://jackyfl.github.io/vitsplit.github.io/.
@@ 1 Introduction
To leverage prior knowledge from VFMs, previous VFM adapters such as ViT-Adapter [14] or ViT-CoMer [78] primarily adopt a two-branch architecture (see Fig. 1(a)). Such a design enables the adapter to integrate low-level features from a convolutional neural network (CNN) with global features from a vision transformer (ViT)-based VFM. While this architecture has demonstrated promising results across various downstream tasks, certain design aspects may affect training efficiency. From Fig. 1(a), we identify two main issues of inefficiency. First, the interaction between the CNN and ViT branches across multiple stages requires gradients to be back-propagated through all layers of the model during training. This results in increased computational and memory costs, which become more significant as the size of the VFM grows. Second, all components need to be tuned during the training process to achieve optimal performance. Specifically, for tasks like segmentation, a large head such as Mask2Former [16] is tuned, and its size is nearly equivalent to that of the VFM backbone.
@@ 2.2 PEFT and VFM adapters
Current VFM adapters (ViT-Adapter [14], ViT-CoMer [78]) aim to enhance full fine-tuning performance by incorporating the inductive bias from the CNN branch with spatial prior. These adapters typically require tuning the whole backbone to achieve optimal performance, resulting in better performance than PEFT methods. The interaction between CNN and ViT features is achieved through cross-attention [14], self-attention [78] or mixed [89] across several layers. In contrast to PEFT methods and existing VFM adapters, our ViT-Split approach keeps the entire backbone frozen, introducing two lightweight heads for separate tuning. This strategy proves to be both efficient and effective across various tasks.
@@ 4 Experiments
We conduct experiments on three tasks—semantic segmentation, object detection, and VQA—using well-established benchmarks such as COCO [51], ADE20K [91], and CityScapes [20], among others. Additionally, we present MDE results in the Appendix. Next, we perform an ablation study to further evaluate ViT-Split’s performance. A uniform selection strategy is applied to all experiments in this section, while results for the sparse gate are provided in the Appendix.
@@ 4.3 Visual Question Answering
Settings. We also present VQA results using the popular visual large language model (VLLM), LLaVA-1.5 [53]. This model comprises a CLIP-L visual encoder for encoding images, an MLP connector for projecting visual tokens into the language space, and a Vicuna-based LLM [17] for generating language tokens. In our modified LLaVA, we replace the original MLP projector with our ViT-Split. To comprehensively evaluate the effectiveness of our ViT-Split, we utilize both academic-task-oriented benchmarks ( VQA-v2 [27], VizWiz [30], SciQA-IMG [60]), and instruction-following LLM benchmarks (POPE [49], MMBench [55], LLaVA-Wild [54], MM-Vet [84]). Following [53], we first pretrain our ViT-Split using 558K image-text pairs, and subsequently fine-tune both ViT-Split and the LLM with 665K mixed data pairs. For more detailed information regarding the hyperparameter settings, please refer to the Appendix.
@@ A.2 Sizes of various heads
A.3 Architecture details of various used VFMs
@@ Appendix C Motivation of freezing the backbone
Appendix D More results
D.1 Hyper-parameter sensitivity analysis
@@
D.2 Visualization
D.2.1 CKA analysis of other VFMs
@@
D.2.2 CKA analysis of different DINOv2 sizes
@@
D.2.3 More layer feature comparison
@@
D.2.4 Semantic segmentation and instance segmentation results
@@
D.3 Training efficiency comparison
@@
D.4 Longer training time
@@
D.5 Monocular depth estimation
@@
D.6 Segmentation on Pascal Context

### v2

@@ Abstract
Vision foundation models (VFMs) have demonstrated remarkable performance across a wide range of downstream tasks. While several VFM adapters have shown promising results by leveraging the prior knowledge of VFMs, we identify two inefficiencies in these approaches. First, the interaction between convolutional neural network (CNN) and VFM backbone triggers early layer gradient backpropagation. Second, existing methods require tuning all components, adding complexity. Besides, these adapters alter VFM features, underutilizing the prior knowledge. To tackle these challenges, we propose a new approach called ViT-Split, based on a key observation: the layers of several VFMs, like DINOv2, can be divided into two components: an extractor for learning low-level features and an adapter for learning task-specific features. Leveraging this insight, we eliminate the CNN branch and introduce two heads, task head and prior head, to the frozen VFM. The task head is designed to learn task-specific features, mitigating the early gradient propagation issue. The prior head is used to leverage the multi-scale prior features from the frozen VFM, reducing tuning parameters and overfitting. Extensive experiments on various tasks (e.g., segmentation, detection, depth estimation, and visual question answering) validate the effectiveness and efficiency of ViT-Split. Specifically, ViT-Split reduces training time up to 4×4$\times$4 × while achieving comparable or even better results on ADE20K, compared to other VFM adapters. Codes are available: https://jackyfl.github.io/vitsplit.github.io/.
@@ 1 Introduction
To leverage prior knowledge from VFMs, previous VFM adapters such as ViT-Adapter [14] or ViT-CoMer [79] primarily adopt a two-branch architecture (see Fig. 1(a)). Such a design enables the adapter to integrate low-level features from a convolutional neural network (CNN) with global features from a vision transformer (ViT)-based VFM. While this architecture has demonstrated promising results across various downstream tasks, certain design aspects may affect training efficiency. From Fig. 1(a), we identify two main issues of inefficiency. First, the interaction between the CNN and ViT branches across multiple stages requires gradients to be back-propagated through all layers of the model during training. This results in increased computational and memory costs as the size of the VFM grows. Second, all components need to be tuned during training to achieve optimal performance. Specifically, for tasks like segmentation, a large head such as Mask2Former [16] is tuned, and its size is nearly equivalent to that of the VFM backbone.
@@ 2.2 PEFT and VFM adapters
Current VFM adapters (ViT-Adapter [14], ViT-CoMer [79]) aim to enhance full fine-tuning performance by incorporating the inductive bias from the CNN branch with spatial prior. These adapters typically require tuning the whole backbone to achieve optimal performance, resulting in better performance than PEFT methods. The interaction between CNN and ViT features is achieved through cross-attention [14], self-attention [79] or mixed [90] across several layers. By contrast, our ViT-Split keeps the entire backbone frozen, introducing two lightweight heads for separate tuning, which is efficient and effective across various tasks.
@@ 4 Experiments
We conduct experiments on three tasks, semantic segmentation, object detection, and VQA, using well-established benchmarks, e.g., COCO [52], ADE20K [92], CityScapes [20], among others. We also present MDE results in the Appendix. Next, we perform ablation studies to further evaluate ViT-Split’s performance. A uniform selection strategy is applied to all experiments in this section, while results for the sparse gate are provided in the Appendix.
@@ 4.3 Visual Question Answering
Settings. We also present VQA results using the popular visual large language model (VLLM) [50], LLaVA-1.5 [54]. This model comprises a CLIP-L visual encoder for encoding images, an MLP connector for projecting visual tokens into the language space, and a Vicuna-based LLM [17] for generating language tokens. In our modified LLaVA, we replace the original MLP projector with our ViT-Split. To comprehensively evaluate the effectiveness of our ViT-Split, we utilize both academic-task-oriented benchmarks ( VQA-v2 [27], VizWiz [30], SciQA-IMG [61]), and instruction-following LLM benchmarks (POPE [49], MMBench [56], LLaVA-Wild [55], MM-Vet [85]). Following [54], we first pretrain our ViT-Split using 558K image-text pairs, and subsequently fine-tune both ViT-Split and the LLM with 665K mixed data pairs. For more detailed information regarding the hyperparameter settings, please refer to the Appendix.
@@ A.2 Sizes of various heads
A.3 Details of tuning the VLLM
LLaVA-1.5 employs a CLIP-based vision encoder for image encoding. We introduce a single-layer task head copied from CLIP’s original final layer (i.e., $K_{t}=1$) and utilize only the last-layer feature of CLIP as the input to the prior head (i.e., $K_{p}=1$). We replace the original MLP projector in LLaVA-1.5 with with our ViT-Split for two-stage training. The training follows the same hyperparameter settings as the original LLaVA-1.5.
A.4 Architecture details of various used VFMs
@@ Appendix C Motivation of freezing the backbone
Appendix D Explanation of the lower performance on detection task
We acknowledge that the performance difference between ViT-Split and ViT-CoMer on Mask R-CNN (Tab. 5) is relatively small. However, ViT-Split uses only 90%–95% of ViT-CoMer’s trainable parameters, already demonstrating clear advantages in training efficiency while maintaining comparable accuracy. The primary reason ViT-Split does not significantly outperform other VFM-adapters lies in the relatively weak task alignment of the prior features from DINOv2 for object detection tasks. Unlike DETR-style models, which are pre-trained with strong detection-oriented objectives, self-supervised models like DINOv2 tend to provide less directly transferable features for detection. This necessitates using more layers in the task head (i.e., larger $K_{t}$), effectively making ViT-Split rely more on fine-tuning, similar to other VFM-adapters. As self-supervised models begin to offer stronger detection-aware priors, we expect ViT-Split to better leverage them and close the gap with current SOTA DETR-style models.
Appendix E More results
E.1 An apple-to-apple comparison with other VFM-adapters on segmentation
We provide an apple-to-apple comparison with the SOTA VFM-adapters in Tab. 14, i.e., ViT-CoMer [79] and ViT-Adapter [14]. All models are trained for 40K iterations on ADE20K, using a UperNet head for the baselines and a linear head for ViT-Split. For VFM-adapters, we adopt a learning rate schedule similar to that used in detection tasks, incorporating layer-wise decay with carefully tuned rates for each baseline to ensure strong performance. Results show that with DINOv2 initialization, ViT-Split consistently outperforms other VFM-adapters across different model sizes. This highlights ViT-Split’s ability to better leverage the strong prior knowledge from DINOv2 without altering the original feature representations, which often results in suboptimal performance in other adapters.
Table 14: VFM-adapter comparison on ADE20K (40K iterations).
Method DINOv2-S DINOv2-B DINOv2-L
mAP #Param mAP #Param mAP #Param
ViT-Adapter 45.4 57.1M 49.9 134.5M 52.2 364.6M
ViT-CoMer 44.6 61.8M 48.4 145.6M 51.8 384.2M
ViT-Split 51.6 10.2M 55.7 40.5M 58.2 88.6M
E.2 Hyper-parameter sensitivity analysis
@@
E.3 Visualization
E.3.1 CKA analysis of other VFMs
@@
E.3.2 CKA analysis of different DINOv2 sizes
@@
E.3.3 More layer feature comparison
@@
E.3.4 Semantic segmentation and instance segmentation results
@@
E.4 Training efficiency comparison
@@
E.5 Longer training time
@@
E.6 Monocular depth estimation
@@
E.7 Segmentation on Pascal Context